OpenStudy (anonymous):

Please help Prove the following (sin(3x)+sin(x))/(sin(3x)-sin(x)=2/(1-tan^2(x))

OpenStudy (anonymous):

So far by using addition identities I got to -2/(2tan2(x)-sec^2(x))

OpenStudy (jdoe0001):

\(\bf \cfrac{sin(3x)+sin(x)}{sin(3x)-sin(x)}=\cfrac{2}{1-tan^2(x)}?\)

myininaya (myininaya):

question: are you allowed to use the following two identities: \[\sin(\alpha)+\sin(\beta)=2 \sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha-\beta}{2}) \\ \sin(\alpha)-\sin(\beta)=2 \cos(\frac{\alpha+\beta}{2})\sin(\frac{\alpha-\beta}{2})\] if not I will go ahead and check by addition identities that I can get to what you have ?

OpenStudy (anonymous):

Yes @jdoe0001 And I am not allowed to use those @myininaya

myininaya (myininaya):

hmmm I'm not getting that expression after using those sum identities but that doesn't mean what you have is wrong let's assume it is right try writing sec^2(x) in terms of tan^2(x) using one of the Pythagorean identities

OpenStudy (anonymous):

Alright I'll try that

myininaya (myininaya):

yep it looks like you were almost there

OpenStudy (anonymous):

Wow I was really close

OpenStudy (anonymous):

Thanks a lot! :)

OpenStudy (anonymous):

But just to make sure we just multiply by -1/-1 at the end to get the one on the RHS right?

myininaya (myininaya):

yep

OpenStudy (anonymous):

k thanks! :)

myininaya (myininaya):

You use that Pythagorean identity I mentioned collect like terms on bottom then last step multiply by -1/-1

myininaya (myininaya):

np